1. |x-2|=8 2. 0=|2+6x| 3. |2x+5|+6=3
How many solutions for each?

Answer by stanbon(75887)

1. |x-2|=8
x-2 = 8 or x-2 = -8
x = 10 or x = -6
Two solutions
-------------------------

2. 0=|2+6x|
2+6x = 0
x = -1/3
One solution
------------------


3. |2x+5|+6=3
|2x+5) = -3
That is impossible
No solutions
=====================
Cheers,
Stan H.
===============